OpenStudy (summersnow8):

pleeeease help me!! tan x=7/24 , x Є [π,3π/2] cos (x/2) =?? sin (x/2) =?? tan (x/2)= ??

OpenStudy (anonymous):

Are you using radians or degrees?

OpenStudy (summersnow8):

im in degree mode, but the answers are suppose to have sqrts in them

OpenStudy (anonymous):

k one sec.

OpenStudy (anonymous):

Is it multiple choice?

OpenStudy (summersnow8):

no, i now the answer, I just dont know how to get it

OpenStudy (summersnow8):

know*

OpenStudy (anonymous):

Can you type to full question? Or is that it?

OpenStudy (summersnow8):

that is the question

OpenStudy (anonymous):

hmm well... I'm gonna need help lol One sec.

OpenStudy (anonymous):

@amistre64 @tcarroll010 Got time to help?

OpenStudy (anonymous):

Or, @Luigi0210 :P

OpenStudy (luigi0210):

Use the half-angle formula?

OpenStudy (anonymous):

I doubt that would work but I don't know...

OpenStudy (luigi0210):

Naw, I think @tcarroll010 would be more helpful than me.

OpenStudy (summersnow8):

answers i was given: sin (x/2) = 7 sqrt(2) / 10 cos (x/2)= sqrt(2) / 10 tan(x/2)= -7

OpenStudy (amistre64):



OpenStudy (luigi0210):

If I remember correctly anything over 2 in a trig function means half-angle..

OpenStudy (summersnow8):

yeah i have tried using a triangle, and the half angle identity, but i cant seem to get the right answers when solving

OpenStudy (anonymous):

Just curious, what class is this for?

OpenStudy (summersnow8):

trig

OpenStudy (anonymous):

Wow... lol I'm only in Algebra 2... Probably why I'm not much help. :P

OpenStudy (anonymous):

I saw cos sin and tan and thought I might be able to do it but I obviously haven't gotten that far. :P

OpenStudy (amistre64):